OpenStudy (anonymous):

What is the slope of the line through the points P(–5, –6) and Q(–12, 5)?

OpenStudy (anonymous):

A. -11/7 B. 17/11 C. -7/11 D. 11/17

OpenStudy (anonymous):

A

OpenStudy (anonymous):

m=(5+6/(-12+5) m=-11/7

OpenStudy (anonymous):

thanks!